# Robust Identity Perceptual Watermark Against Deepfake Face Swapping

### v2

@@ Abstract
Notwithstanding offering convenience and entertainment to society, Deepfake face swapping has caused critical privacy issues with the rapid development of deep generative models. Due to imperceptible artifacts in high-quality synthetic images, passive detection models against face swapping in recent years usually suffer performance damping regarding the generalizability issue. Therefore, several studies have been attempted to proactively protect the original images against malicious manipulations by inserting invisible signals in advance. However, the existing proactive defense approaches demonstrate unsatisfactory results with respect to visual quality, detection accuracy, and source tracing ability. In this study, to fulfill the research gap, we propose the first robust identity perceptual watermarking framework that concurrently performs detection and source tracing against Deepfake face swapping proactively. We assign identity semantics regarding the image contents to the watermarks and devise an unpredictable and nonreversible chaotic encryption system to ensure watermark confidentiality. The watermarks are encoded and recovered by jointly training an encoder-decoder framework along with adversarial image manipulations. Falsification and source tracing are accomplished by justifying the consistency between the content-matched identity perceptual watermark and the recovered robust watermark from the image. Extensive experiments demonstrate state-of-the-art detection performance on Deepfake face swapping under both cross-dataset and cross-manipulation settings.
1 Introduction
@@
Most existing countermoves [38, 22, 60, 46, 13, 33, 50, 35] regarding Deepfake face swapping follow the pipeline of devising a deep detection model that passively investigates and analyzes the underlying synthetic artifacts to distinguish the fake images. While accomplishing progresses in the early stages, passive detection suffers the generalization challenge such that obvious performance damping occurs when attempting to detect fake materials produced by unknown synthetic algorithms [45]. Moreover, following rapid improvements in deep generative algorithms, it has become burdensome to locate perceivable manipulation traces within high-quality synthetic images. To overcome this issue, several recent studies consider executing proactive defense against synthetic manipulations. In other words, the goal is to protect the original images by inserting imperceptible signals in advance of posting in public and being manipulated by Deepfake unexpectedly.
@@
The proactive approaches can be categorized into two main strategies, namely, distorting [15, 44, 48] and watermarking [55, 31, 3, 30, 25]. The former learns a distortion that nullifies potential synthetic manipulations by inserting it into the original images, and the latter encodes semi-fragile watermarks into original images and determines the authenticity based on the existence of watermarks afterward. Both ways aim to proactively protect the original images without affecting the visual qualities. While demonstrating preliminary initiations, the existing approaches suffer several flaws. Firstly, in the current research stage, both the distorting and watermarking approaches lack generalization ability on unseen datasets when facing unknown synthetic manipulations. Secondly, although the semi-fragile watermarks can help determine real and fake based on their existence, they are prone to be unstable when facing unavoidable common image processing operations. Moreover, watermarks that are fragile to malicious synthetic manipulations cannot help conduct source tracing on the victim target images simultaneously. In particular, in an integral chain of evidence for the forensic investigation in the cybercrime cases related to Deepfake face swapping, besides correctly addressing the falsification, it is crucial to retrieve intelligence information of the source materials [1, 20, 51]. In other words, tracing the original target images along with the detection task can help fulfill the evidence integrality. Lastly, while preventing privacy issues, directly adding distortions not only unavoidably degrades the visual qualities of the images [11], but also unfavorably disables benign adoptions of Deepfake face swapping.
@@
To address the aforementioned issues, in this study, we propose a novel robust identity perceptual watermarking framework that promotes proactive defense against Deepfake face swapping. In general, we establish semantically meaningful identity perceptual watermarks that are robust and invisible based on image contents and embed them into original images for detection and source tracing purposes simultaneously, while allowing benign usages of face swapping algorithms. First, based on the identities of facial images, we construct securely protected identity perceptual watermarks by introducing a chaotic encryption system. Then, we devise an end-to-end robust watermarking framework that consists of an encoder-decoder architecture to be trained with common and face swapping manipulation pools. Relying on the promising watermark recovery performance, the authenticity of a candidate image is determined by the consistency between the content-matched identity perceptual watermark and the robustly recovered watermark from the image. The source tracing is simultaneously achieved with respect to the identity semantics carried by the recovered watermark. Extensive experiments prove the generalization ability of our proposed approach under cross-dataset and cross-manipulation settings with average watermark recovery accuracies all above 96%, strongly outperforming the contrastive methods. The outstanding detection performance against Deepfake face swapping with AUC scores all above 97% also consistently outperforms the state-of-the-art passive Deepfake detectors.
@@
  • We propose a novel idea of identity perceptual watermarks based on image contents to proactively defend against Deepfake face swapping regarding content-watermark consistencies. A chaotic encryption system is collaboratively devised to ensure watermark confidentiality.
  • We present an encoder-decoder framework to robustly and invisibly embed the watermarks into facial images. To the best of our knowledge, we are the first to concurrently achieve detection and source tracing for Deepfake face swapping by embedding a single watermark.
  • Experiments qualitatively and quantitatively demonstrate the promising watermark recovery accuracy and Deepfake face swapping detection performance of our method along with outstanding generalization ability, outperforming the related state-of-the-art algorithms.
2 Related Work
Existing forensic investigations on Deepfake face swapping mostly rely on devising passive detectors [38, 22, 60, 13, 32, 33, 50] that analyze traces of suspected images in various feature domains. In view of the rising challenge prompted by the improving synthetic quality, instead of searching for synthetic artifacts, proactive defense aims to protect original images by inserting imperceptible signals before potential manipulations. In other words, while malicious face swapping happens after the target image is publicly posted (e.g., on Twitter or Instagram), proactive actions are performed in advance in two ways: distortions and watermarks.
2.1 Proactive Distortions
Proactive distortions [37, 17, 57, 56, 2] are usually inserted into pristine images to nullify the synthetic models following $\tilde{x}=x+\eta$, where $x$ and $\tilde{x}$ are the pristine images before and after the learned invisible distortion $\eta$ is added, respectively. In particular, recently, Anti-Forgery [44] exploits the vulnerability of GAN-based models and supervisely learns a particular distortion in the lab color space for a specific pristine image regarding each synthetic algorithm. CMUA [15] presents a two-level perturbation fusion strategy and heuristically learns a universal distortion that can disrupt multiple synthetic algorithms. TAFIM [2] introduces image-specific perturbations with better robustness via attention-based fusion and refinement. Besides adding invisible perturbations that could heavily distort the synthetic results, Wang et al. [48] further promoted a detector that distinguishes the distorted images with high accuracy since machines do not share similar perspectives as human eyes.
2.2 Proactive Watermarks
@@
Figure 1: Framework of the proposed approach. An original clean image $I$ is first passed to generator $G_{m}$ to generate the corresponding identity perceptual watermark $m$. Then, $I$ and $m$ are fed to the encoder for watermark embedding. Manipulation pools $P_{\textrm{common}}$ and $P_{\textrm{swap}}$ help preserve watermark robustness during supervision. A discriminator helps maintain the visual quality. In the end, the watermark $m_{\textrm{rec}}$ is recovered by passing the manipulated image to the decoder.
@@
Regarding the relatively novel term of proactive defense, few approaches have been attempted with watermarks. Yu et al. [58, 59] assigned particular fingerprints to images via robust watermarking regarding generative models for the attribution purpose. Different from the robust watermarking studies, FaceGuard [55] and FaceSigns [31] propose semi-fragile watermarks such that they stay robust to benign image post-processing operations but become fragile when facing malicious face manipulations. The authors claimed to distinguish fake materials if no watermark is detected in candidate images. Recently, a separable watermarking framework, SepMark [51], is conducted with two watermark decoders, a Tracer and a Detector, where the former recovers robust watermarks for source tracing, and the latter extracts semi-fragile watermarks to determine real and fake. To the best of our knowledge, robust watermarking has been barely attempted to detect Deepfake face swapping adaptively, and performing detection and source tracing simultaneously with a single watermark remains largely unexplored in the current research domain.
3 Methodology
3.1 Overview
@@
In this paper, we propose an identity perceptual watermarking framework to adaptively embed imperceptible robust watermarks into images and proactively achieve detection and source tracing against Deepfake face swapping by precisely recovering the encoded watermarks. The model framework is demonstrated in Fig. 1. First, we devise a watermark generator $G_{m}$ to securely prepare the binary watermarks by assigning identity perceptual semantics with respect to the image contents. Then, the encoder is fed with the clean image $I$ and the corresponding watermark $m$ for encoding. To guarantee robustness, we apply a common manipulation pool $P_{\textrm{common}}$ and a face swapping manipulation pool $P_{\textrm{swap}}$ as adversaries to the watermarked image $I_{\textrm{rec}}$ before recovering the watermark $m_{\textrm{rec}}$ via the decoder. Meanwhile, a discriminator that distinguishes the watermarked and clean images is adopted in the training stage to enhance image visual quality. By assigning constraints as discussed in Section 3.5, the framework is trained to satisfy our objectives.
@@
In practice, for an original image $I$ with facial identity $t$, the identity perceptual watermark $m$ is produced by $G_{m}$ and embedded into $I$ to generate the watermarked image $I_{\textrm{rec}}$ via the encoder. Upon the malicious face swapping manipulation, a synthetic facial image $I_{\textrm{fake}}$ is derived by swapping identity $s$ onto $I_{\textrm{rec}}$. By utilizing the decoder, the watermark $m_{\textrm{rec}}$ which is faithfully consistent with $m$ can be robustly recovered from $I_{\textrm{fake}}$. Since $I_{\textrm{fake}}$ contains facial identity $s$, its corresponding watermark $m_{s}$ can be generated using $G_{m}$, and demonstrating the inconsistency between $m_{\textrm{rec}}$ and $m_{s}$ can accomplish the detection task on Deepfake face swapping. On the other hand, as $m_{\textrm{rec}}$ is robustly recovered, the trusted original image $I$ with identity perceptual watermark $m$ can be traced by inspecting the matchings between identities and watermarks in an officially recorded hash map.
3.2 Identity Perceptual Watermarks
@@
Unlike early frameworks that enforce random initialization, we derive the watermark for each facial image by analyzing the identity semantics. In particular, we propose an identity perceptual watermark generator $G_{m}$ that generates binary watermark $m$ with length $l$ based on the facial identity of image $I$. Inspired by hashing algorithms [36, 24], the pipeline of $G_{m}$ follows two critical characteristics: collision free and one-way computational.
@@
Collision Free. Since the goal is to detect Deepfake face swapping by addressing inconsistencies between the recovered and content-matched watermarks of candidate images, it is necessary to ensure that no two identical watermarks belong to different facial identities. Face recognition, a task to identify different facial identities by producing distinguishable embeddings, has been proved the ability on various large datasets with reliable models. In this study, impressed by the outstanding achievements, we adopt well-performed face recognition algorithms that retrieve identity embeddings in the form of high-dimensional vectors. To match up the required watermark length $l$, we first select a large dataset that contains sufficient unique identities and retrieve the identity embeddings $E$ with dimension $d_{E}$ accordingly. Then, we apply the principal component analysis (PCA) [34] to $E$ and learn a transform rule $\phi$ that projects the identity embeddings of the images from dimension $d_{E}$ down to $l$ while optimally preserving the original correlations between facial identities with respect to the principal components. Thereafter, for the purpose of constructing binary watermarks, we conduct a unit-variance scaling on the projected embeddings $E_{\textrm{pca}}$ to the range between 0 and 1 by
@@
where $E_{\textrm{scale}}$ denotes the scaled embeddings with respect to the standard deviation, which favorably preserves the distribution of $E_{\textrm{pca}}$. In the end, we set a cutoff value $0\leq c\leq 1$ to round the entries of $E_{\textrm{scale}}$ to binary values accordingly and obtain the identity perceptual watermarks $E_{\textrm{binary}}$. Besides, a watermark collision check is arranged in Section 4.4 to validate the collision-free characteristic.
One-Way Computational. Since we concurrently conduct detection and source tracing, an adversary with sufficient knowledge of our work may aim to replace the encoded watermark with a content-matched one. To prevent hostile attacks [42, 6], we enforce the watermark to be one-way computational, that is, it is insufficient to recover the watermark construction pipeline by merely acquiring the watermark from the image. Specifically, besides keeping $\phi$ encapsulated, we design a chaotic encryption [29] system on $E_{\textrm{binary}}$ using a logistic map
@@
where $p$ and $q$ are pre-defined prime constants. Finally, the encrypted identity perceptual watermark $m$ for image $I$ is derived by a logical exclusive OR (XOR) operation between $K$ and the unencrypted watermark from $E_{\textrm{binary}}$. Relying on the sensitive dependence on the initial condition and chaotic map nonlinearity characteristics, the final watermarks leveraged in the watermarking framework are unpredictable and nonreversible unless having access to Eqn.(2) and Eqn.(3) including the critical constant coefficients. Specifically, the entry values in $X$ vary drastically with respect to the values of $x_{0}$ and $r$, and the watermarks are thus securely protected. On the other hand, with authority on $K$, performing an extra XOR operation on $m$ using $K$ can easily decrypt the watermarks back to the ones in $E_{\textrm{scale}}$ to trace back for the identity information.
Figure 3: Visual effects of Deepfake face swapping image manipulations on the watermarked images. The first column refers to the source images that provide facial identity for face swapping. Columns 2 and 3 present the raw and watermarked target images. Every two columns of the remaining exhibit the face swapping visualizations on the raw and watermarked target images.
3.3 Watermark Encoding
For an input image $I$, we insert the watermark $m$ using an encoder that is mainly based on convolutional neural networks (CNN) with squeeze-and-excitation networks (SENet) [12] and diffusion blocks. Specifically, the image features and watermark features are separately analyzed and then concatenated to reconstruct the watermarked image $I_{\textrm{rec}}$.

### v3

@@ Abstract
Notwithstanding offering convenience and entertainment to society, Deepfake face swapping has caused critical privacy issues with the rapid development of deep generative models. Due to imperceptible artifacts in high-quality synthetic images, passive detection models against face swapping in recent years usually suffer performance damping regarding the generalizability issue in cross-domain scenarios. Therefore, several studies have been attempted to proactively protect the original images against malicious manipulations by inserting invisible signals in advance. However, existing proactive defense approaches demonstrate unsatisfactory results with respect to visual quality, detection accuracy, and source tracing ability. In this study, to fulfill the research gap, we propose a robust identity perceptual watermarking framework that concurrently performs detection and source tracing against Deepfake face swapping proactively. We innovatively assign identity semantics regarding the image contents to the watermarks and devise an unpredictable and nonreversible chaotic encryption system to ensure watermark confidentiality. The watermarks are robustly encoded and recovered by jointly training an encoder-decoder framework along with adversarial image manipulations. For a suspect image, falsification is accomplished by justifying the consistency between the content-matched identity perceptual watermark and the recovered robust watermark, without requiring the ground-truth. Moreover, source tracing can be accomplished based on the identity semantics that the recovered watermark carries. Extensive experiments demonstrate state-of-the-art detection and source tracing performance against Deepfake face swapping with promising watermark robustness for both cross-dataset and cross-manipulation settings.
I Introduction
@@
Most existing countermoves [2, 3, 4, 5, 6, 7, 8, 9] regarding Deepfake face swapping follow the pipeline of devising a deep detection model that passively investigates and analyzes the underlying synthetic artifacts to distinguish the fake images. While accomplishing progresses in the early stages, passive detection suffers the generalization challenge such that obvious performance damping occurs when attempting to detect fake materials produced by unknown synthetic algorithms [10]. Moreover, following rapid improvements in deep generative algorithms, it has become burdensome to locate perceivable manipulation traces within high-quality synthetic images. To overcome this issue, several recent studies consider executing proactive defense against synthetic manipulations. In other words, the goal is to protect the original images by inserting imperceptible signals in advance of posting to the public and being manipulated by Deepfake unexpectedly.
@@
The proactive approaches against Deepfake face swapping can be categorized into two main strategies, namely, distorting [11, 12, 13] and watermarking [14, 15, 16, 17, 18]. The former learns a distortion that nullifies potential synthetic manipulations by inserting it into the original images, and the latter encodes watermarks into original images and determines the authenticity based on the recovered watermarks afterward. Both ways aim to proactively protect the original images without affecting the visual qualities. While demonstrating preliminary initiations, the existing approaches suffer several flaws. Firstly, in the current research stage, both the distorting and watermarking approaches lack generalization ability on unseen datasets when facing unknown synthetic manipulations. Secondly, although the semi-fragile watermarks can help determine real and fake based on their existence, they are prone to be unstable when facing unavoidable common image processing operations. Moreover, watermarks that are fragile to malicious synthetic manipulations cannot help conduct source tracing on the victim target images simultaneously. In particular, in an integral chain of evidence for the forensic investigation in the cybercrime cases related to Deepfake face swapping, besides correctly addressing the falsification, it is crucial to retrieve intelligence information of the source materials [19, 20, 21]. In other words, tracing the original target images along with the detection task can help fulfill the evidence integrality. On the other hand, although few robust watermarking studies have been attempted, they generally determine real and fake relying on the ground-truths because the watermarks themselves are semantically void, making the proactive detection pipeline incomplete when the ground-truths are unavailable in real-life cases. Lastly, while preventing privacy issues, directly adding distortions not only unavoidably degrades the visual qualities of the images [22], but also unfavorably disables benign adoptions of Deepfake face swapping.
@@
To address the aforementioned issues, in this study, we propose a novel robust identity perceptual watermarking framework that promotes proactive defense against Deepfake face swapping. In general, we establish semantically meaningful identity perceptual watermarks that are robust and invisible based on image contents and embed them into original images for detection and source tracing purposes simultaneously, while allowing benign usages of face swapping algorithms. First, based on the identities of facial images, we construct securely protected identity perceptual watermarks by introducing a chaotic encryption system. Then, we devise an end-to-end robust watermarking workflow that consists of an encoder-decoder architecture to be trained with common and face swapping manipulation pools. The model is pre-trained with the common image processing operations to accomplish robustness with good visual quality when facing benign manipulations. Thereafter, we tune it with a single representative face swapping algorithm to achieve the objectives. Relying on the promising watermark recovery performance, the authenticity of a candidate image is determined by evaluating the consistency between its content-matched identity perceptual watermark and the robustly recovered watermark, without requiring ground-truth watermarks. The source tracing is simultaneously achieved with respect to the identity semantics carried by the recovered watermark. Extensive experiments prove the generalization ability of our proposed approach under cross-dataset and cross-manipulation settings, with average watermark recovery accuracies strongly outperforming the contrastive methods. Further, the outstanding detection performance against four GAN-based and four diffusion-based face swapping algorithms also consistently outperforms the state-of-the-art passive Deepfake detectors. Meanwhile, the average top-5, top-3, and top-1 source tracing accuracies of $98.10\%$, $97.72\%$, and $96.56\%$ against the eight face swapping models demonstrate the reliable ability of our proposed method to trace back to the original target facial identities.
@@
  • We propose a novel idea of identity perceptual watermarks based on facial image contents to proactively defend against Deepfake face swapping regarding content-watermark consistencies, without requiring ground-truths. A chaotic encryption system is collaboratively devised to ensure watermark confidentiality.
  • We present an encoder-decoder workflow to invisibly embed watermarks into facial images and robustly recover them. Upon maintaining watermark robustness, our approach can concurrently achieve detection and source tracing for Deepfake face swapping by embedding and decoding a single robust watermark.
  • Experiments qualitatively and quantitatively demonstrate the promising watermark recovery accuracy, Deepfake face swapping detection performance, and source tracing ability of our proposed approach, along with outstanding generalization ability against both GAN-based and diffusion-based generative models, outperforming state-of-the-art algorithms.
II Related Work
Existing forensic investigations on Deepfake face swapping mostly rely on devising passive detectors [2, 3, 4, 6, 23, 7, 8] that analyze traces of suspected images in various feature domains. In view of the rising challenge prompted by the improving synthetic quality due to well-developed generative models, instead of searching for synthetic artifacts, proactive defense aims to protect original images by inserting imperceptible signals before potential manipulations. In other words, while malicious face swapping happens after the target image is publicly posted (e.g., on Twitter or Instagram), proactive actions are usually performed in advance in two ways: distortions and watermarks.
II-A Proactive Distortions
Proactive distortions [24, 25, 26, 27, 28] are usually inserted into pristine images to nullify the synthetic models following $\tilde{x}=x+\eta$, where $x$ and $\tilde{x}$ are the pristine images before and after the learned invisible distortion $\eta$ is added, respectively. An early work [29] introduces the oscillating GAN (OGAN) attack, which successfully creates visible oscillations in synthetic facial images. Recently, Anti-Forgery [12] exploits the vulnerability of GAN-based models and supervisely learns a particular distortion in the lab color space for a specific pristine image regarding each synthetic algorithm. CMUA [11] presents a two-level perturbation fusion strategy and heuristically learns a universal distortion that can disrupt multiple synthetic algorithms. TAFIM [28] introduces image-specific perturbations with better robustness via attention-based fusion and refinement. Besides adding invisible perturbations that could heavily distort the synthetic results, Wang et al. [13] further promoted a detector that distinguishes the distorted images with high accuracy since machines do not share similar perspectives as human eyes.
II-B Proactive Watermarks
@@
Figure 1: Framework of the proposed approach. An original clean image $I$ is first passed to generator $G_{m}$ to generate the corresponding identity perceptual watermark $m$. Then, II and $m$ are fed to the encoder for watermark embedding, which derives $I_{\textrm{rec}}$. Manipulation pools $P_{\textrm{common}}$ and $P_{\textrm{swap}}$ containing common and Deepfake manipulations help enhance watermark robustness during supervision. Besides, a discriminator helps maintain the visual quality. In the end, the watermark $m_{\textrm{rec}}$ is recovered by passing the manipulated image to the decoder.
@@
Regarding the relatively novel term of proactive defense in digital forensics, few approaches have been attempted with watermarks [40, 21, 41, 42, 43]. Specifically, Yu et al. [40, 44] assigned particular fingerprints to images via robust watermarking regarding generative models for the attribution purpose. Wang et al. [45] raised the importance of Deepfake provenance tracking and designed FakeTagger to perform image tagging via a novel channel-coding injection operation. Different from the robust watermarking studies, FaceGuard [14] and FaceSigns [15] propose semi-fragile watermarks such that they stay robust to benign image post-processing operations but become fragile when facing malicious face manipulations. The authors claimed to distinguish fake materials if no watermark is detected in candidate images. Meanwhile, Zhao et al. [41] introduced a semi-fragile watermark that performs promisingly against the classical attacks including copy-move, inpainting, and splicing. Later, IMD [46] and MaLP [47] are proposed to detect Deepfake faces by inserting pre-learned semi-fragile templates and computing the cosine similarity between extracted and ground-truth templates, where MaLP is further expanded to all image contents and perform localization based on IMD. Recently, a separable watermarking framework, SepMark [21], is conducted with two watermark decoders, a Tracer and a Detector, where the former recovers robust watermarks for source tracing, and the latter extracts semi-fragile watermarks to determine real and fake. Similarly, FakeTracer [48, 49] introduces robust and semi-fragile traces, STrace and ETrace, to detect Deepfake face swapping according to the co-existence of the traces. Differently, FakeTracer requires the traces to be implanted into images before the training phase of face swapping models such that the generative models learn to maintain only STrace during inference. To the best of our knowledge, compared to the abundance in defending against facial attribute editing, robust watermarking has been barely attempted to detect Deepfake face swapping adaptively. Moreover, performing detection and source tracing simultaneously without ground-truths via a single watermark remains largely unexplored in the current research domain.
III Methodology
III-A Overview
@@
In this paper, we propose a novel robust identity perceptual watermarking framework to adaptively embed imperceptible robust watermarks into images and proactively achieve detection and source tracing against Deepfake face swapping by precisely recovering the encoded watermarks. The framework of our approach is demonstrated in Fig. 1. First, we devise a watermark generator $G_{m}$ to securely prepare the binary watermarks by assigning identity perceptual semantics with respect to the image contents. Then, the encoder is fed with the clean image II and the corresponding identity perceptual watermark $m$ for encoding. To guarantee robustness, we apply a common manipulation pool $P_{\textrm{common}}$ and a face swapping manipulation pool $P_{\textrm{swap}}$ as adversaries to the watermarked image $I_{\textrm{rec}}$ before recovering the watermark $m_{\textrm{rec}}$ via the decoder. Meanwhile, a discriminator that distinguishes the watermarked and clean images is adopted in the training stage to enhance image visual quality. By assigning constraints as discussed in Section III-E, the framework is trained to satisfy our objectives of watermark robustness.
@@
In practice, for an original image $I$ with facial identity $t$, the identity perceptual watermark $m$ is produced by $G_{m}$ and embedded into $I$ to generate the watermarked image $I_{\textrm{rec}}$ via the encoder. Upon malicious face swapping manipulations, a synthetic facial image is derived by swapping a different facial identity onto $I_{\textrm{rec}}$. By utilizing the decoder, the watermark $m_{\textrm{rec}}$, which is faithfully consistent with $m$, can be robustly recovered from the synthetic facial image. Then, the identity perceptual watermark of the synthetic facial image can be obtained via $G_{m}$. Since it contains a different facial identity, its corresponding watermark is different from the robustly recovered watermark $m_{\textrm{rec}}$, and demonstrating the inconsistency between the two watermarks accomplishes Deepfake detection on the synthetic facial image. On the other hand, if the watermark-protected image $I_{\textrm{rec}}$ does not undergo face swapping but certain common image operations, the identity perceptual watermark obtained via $G_{m}$ from the suspect image is identical to $m_{\textrm{rec}}$, which is then classified as real content. Moreover, as $m_{\textrm{rec}}$ is robustly recovered, the original image $I$ with identity perceptual watermark $m$ can be traced based on watermark semantics by inspecting the matching between identities and watermarks.
III-B Identity Perceptual Watermarks
@@
Unlike early approaches that enforce random watermark initialization, we derive the watermark for each facial image regarding the corresponding identity semantics, which then allows the absence of ground-truths in proactive Deepfake detection. In particular, we propose an identity perceptual watermark generator $G_{m}$ that derives a securely encrypted binary watermark $m$ of length $l$ based on the facial identity of the image $I$. Specifically, inspired by hashing algorithms [50, 51], the pipeline of $G_{m}$ follows two critical characteristics: collision free and one-way computational.
@@
Collision Free. Since the goal is to detect Deepfake face swapping by addressing inconsistencies between the recovered and content-perceptual watermarks of candidate suspect images, it is necessary to ensure that no two identical watermarks belong to different facial identities. Face recognition, a task to identify different facial identities by producing distinguishable embeddings, has been proved the ability on various large facial datasets with reliable models. In this study, impressed by the outstanding achievements, we adopt well-performed face recognition algorithms that retrieve identity embeddings in the form of high-dimensional vectors. To match up the required watermark length $l$, we first select a large dataset that contains sufficient unique identities and retrieve the identity embeddings $E$ with dimension $d_{E}$ accordingly. Then, we apply the principal component analysis (PCA) [52] to $E$ and learn a transform rule $\phi$ that projects the identity embeddings of the images from dimension $d_{E}$ down to $l$ while optimally preserving the original correlations between facial identities with respect to the principal components. Thereafter, for the purpose of constructing binary watermarks, we conduct a unit-variance scaling on the projected embeddings $E_{\textrm{pca}}$ to the range between 0 and 1 by
@@
where $E_{\textrm{scale}}$ denotes the scaled embeddings with respect to the standard deviation, which favorably preserves the distribution of $E_{\textrm{pca}}$. In the end, we set a cutoff value $0\leq c\leq 1$ to round the entries of $E_{\textrm{scale}}$ to binary values accordingly, and obtain the identity perceptual watermarks $E_{\textrm{binary}}$. Besides, a watermark collision check is arranged in Section IV-D to validate the collision-free characteristic.
One-Way Computational. Since we concurrently conduct detection and source tracing based solely on robust watermarks, full access to the watermarks is provided to the users. Therefore, an adversary with sufficient knowledge of our work may aim to attack accordingly, such as replacing the encoded watermark with a content-matched one. To prevent hostile attacks [53, 54], we enforce the watermark to be one-way computational, that is, it is insufficient to recover the watermark construction pipeline by merely acquiring the watermark from the image. Specifically, besides keeping $\phi$ encapsulated, we design a chaotic encryption [55] system on $E_{\textrm{binary}}$ using a logistic map
@@
where $p$ and $q$ are pre-defined prime constants. Finally, the encrypted identity perceptual watermark $m$ for image $I$ is derived by a logical exclusive OR (XOR) operation between $K$ and the unencrypted watermark $e_{\textrm{binary}}\in E_{\textrm{binary}}$. Relying on the sensitive dependence on the initial condition and chaotic map nonlinearity characteristics, the final watermarks leveraged in the watermarking framework are unpredictable and nonreversible unless having access to Eqn.(2) and Eqn.(3) including the critical constant coefficients. Specifically, the entry values in $X$ vary drastically with respect to the values of $x_{0}$ and $r$, and the watermarks are thus securely protected. On the other hand, with authority on $K$, performing an extra XOR operation on $m$ using $K$ can easily decrypt the watermarks back to the ones in $E_{\textrm{scale}}$ to trace back for the identity information.
III-C Watermark Encoding
For an input image $I$, we embed the watermark $m$ using an encoder that is mainly based on convolutional neural networks (CNN) with squeeze-and-excitation networks (SENet) [56] and diffusion blocks. Specifically, image and watermark features are separately analyzed and then concatenated to reconstruct the watermarked image $I_{\textrm{rec}}$.